I want to find out how to solve this question:
"How many liters of a 10% solution of acid should mixed with 60% solution of acid to obtain a 20 liters of 50% solution?"
Do I need to use 2 variables? And will my set up of the equation be "x(.10)+y(.60)=20(.50)" ? or are there errors that must be corrected?

Answer by Alan3354(69443)

"How many liters of a 10% solution of acid should mixed with 60% solution of acid to obtain a 20 liters of 50% solution?"
Do I need to use 2 variables? And will my set up of the equation be
"x(.10)+y(.60)=20(.50)" ? or are there errors that must be corrected?
--------------
x(.10)+y(.60)=20(.50)
That's one equation, for the amount of acid.
x + y = 20 is the other.